1.A town has a population of 12000 and grows at 3.5% every year. What will be the population after 7 years, to the nearest whole number?



1
Expert's answer
2021-01-26T03:28:00-0500

"\text{let }p_0 = 12000\text{ city population}"

"\text{then the population of the city in 1 year}"

"p_1= p_0+p_0*\frac{3.5}{100}= p_0*\frac{103.5}{100}=p_0*1.035"

"\text{then the population of the city in 2 year}"

"p_2 = p_1*1.035=p_0*(1.035)^2"

"\text{then the population of the city in n year}"

"p_n = p_0*(1.035)^n"

"\text{then the population of the city in 7 year}"

"p_7 = p_0*(1.035)^7= 12000*(1.035)^7\approx15267"

Answer:15267